OpenStudy (anonymous):

(cos Θ + cos Θ)2 + (cos Θ + cos Θ)2 sin2 Θ 4 8cos2 Θ cos2 Θ

OpenStudy (anonymous):

and now... trigonometry !

OpenStudy (anonymous):

Doesnt it just suck. LOL

OpenStudy (anonymous):

I Think its A.

OpenStudy (anonymous):

yeah for sure

OpenStudy (anonymous):

bruh I hate trig and that's what we're doing in precalc now >.<

OpenStudy (anonymous):

first off no one writes \((\cos(x)+\cos(x))^2\) when it should be \((2\cos(x))^2=4\cos^2(x)\)

OpenStudy (anonymous):

is one of your choices \[8\cos^2(x)\]??

OpenStudy (anonymous):

OH. I Took Pre-calc my junior year. EW.

OpenStudy (anonymous):

Oh, So its C? ;o

OpenStudy (anonymous):

guess so yeah